OpenStudy (anonymous):

Find the following product: (y^2+4y-7)(y+3)

OpenStudy (diyadiya):

\[=y^3+4y^2-7y+3y^2+12y-21\]\[=y^3+7y^2+5y-21\]

jimthompson5910 (jim_thompson5910):

(y^2+4y-7)(y+3) y^2(y+3)+4y(y+3)-7(y+3) (y^2)(y)+(y^2)(3)+(4y)(y)+(4y)(3)+(-7)(y)+(-7)(3) y^3+3y^2+4y^2+12y-7y-21 y^3+7y^2+5y-21 So (y^2+4y-7)(y+3)=y^3+7y^2+5y-21